OpenStudy (anonymous):

if 9tan(theta)=x what are sin(theta) and cos(theta) in terms of 9 and x?

OpenStudy (abb0t):

\(\sf \color{}{tan(\theta )=\frac{x}{9}}\) when tangent = \(\sf \color{red}{\frac{opppsite}{adjacent}}\)

OpenStudy (anonymous):

\[9\tan \theta=x\] \[\tan \theta=\frac{ x }{ 9 }\] \[\tan \theta=\frac{ \sin \theta }{ \cos \theta }=\frac{ x }{ 9 }\]

OpenStudy (anonymous):

 calculate sin theta and cos theta